OpenStudy (anonymous):

Two squares have corresponding sides labeled x² - 7x and 18. What values of x make the squares congruent? @MathematicsTwo squares have corresponding sides labeled x² - 7x and 18. What values of x make the squares congruent? @Mathematics

OpenStudy (anonymous):

x² - 7x = 18 x² - 7x - 18 = 0 (x+2)(x-9)=0 x+2=0 x=-2 x-9=0 x=9 -2 or 9

OpenStudy (anonymous):

Thankk uuuu ur awesomee

OpenStudy (anonymous):

np

OpenStudy (anonymous):

to tell u the truth, i just googled it the most effective way, sort of like cheating

OpenStudy (anonymous):

Oh ok